### v1

# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

Abstract
Prompt optimization (PO) offers a practical alternative to fine-tuning large language models (LLMs), enabling performance improvements without altering model weights. Existing methods typically rely on advanced, large-scale LLMs like GPT-4 to generate optimized prompts. However, due to limited downward compatibility, verbose, instruction-heavy prompts from advanced LLMs can overwhelm lightweight inference models and degrade response quality. In this work, we rethink prompt optimization through the lens of interpretable design. We first identify a set of model-agnostic prompt quality merits and empirically validate their effectiveness in enhancing prompt and response quality. We then introduce MePO, a merit-guided, lightweight, and locally deployable prompt optimizer trained on our preference dataset built from merit-aligned prompts generated by a lightweight LLM. Unlike prior work, MePO avoids online optimization reliance, reduces cost and privacy concerns, and—by learning clear, interpretable merits—generalizes effectively to both large-scale and lightweight inference models. Experiments demonstrate that MePO achieves better results across diverse tasks and model types, offering a scalable and robust solution for real-world deployment [^1] Our model and dataset are available at https://github.com/MidiyaZhu/MePO.
Rethinking Prompt Optimizers: From Prompt Merits to Optimization
Zixiao Zhu1  Hanzhang Zhou1  Zijian Feng1  Tianjiao Li1 Chua Jia Jim Deryl2Mak Lee Onn2Gee Wah Ng2Kezhi Mao1thanks: Corresponding author 1School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 2Home Team Science and Technology Agency, Singapore {zixiao.zhu, hanzhang.zhou, zijian.feng, tianjiao.li, ekzmao}@ntu.edu.sg {Deryl_CHUA, MAK_Lee_Onn, NG_Gee_Wah}@htx.gov.sg
1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks (Achiam et al., 2023; Touvron et al., 2023). However, their output quality remains highly sensitive to prompt phrasing, making effective prompt design crucial for achieving good responses. This has led to growing interest in prompt optimization (PO) (Liu et al., 2023), which aims to improve model performance by refining the input prompt rather than modifying model parameters (Yang et al., ; Zhu et al., 2023).
(a) Prompts from stronger optimization models can degrade the performance of smaller inference models.
(b) Lightweight LLMs fail under complex prompt optimization algorithms.
(c) Lightweight LLMs perform well as PO models with merit guidance.
Figure 1: Empirical analysis of prompt optimization behavior across model scales and optimization algorithms. (a) Raw and GPT-4 optimized prompts are drawn from the Alpaca dataset and Lu et al. (2025). (b) Basic and EvoPrompt examples, along with the optimization algorithm used, are adapted from Guo et al. (2024). (c) The optimized prompt is generated using our merit-guided instruction (Fig. 14). ‘DS’ refers to DeepSeek-R1. Further details are provided in Appx.A.
A dominant PO paradigm is discrete automatic prompt optimization, which employs an advanced, large-scale LLM [^2] In this work, we categorize LLMs like GPT-4 or 70B models as large-scale, while those with fewer than 13B are denoted as lightweight, following ChunLiu et al. (2024). as the prompt optimization model to generate optimized prompts. These prompts are then used as inputs to either similarly sized (Xiang et al., 2025; Ye et al., 2024) or smaller (Guo et al., 2024) inference models for response generation. One line of work adopts online optimization, using advanced API-based models like GPT-4 (Wang et al., ; Zhou et al., 2022; Pryzant et al., 2023). To reduce inference costs, most works typically produce a single optimized task-specific meta-prompt, inserted before or after each query, while leaving the query itself unoptimized (Xiang et al., 2025; Guo et al., 2024; Ye et al., 2024). However, this strategy requires prior knowledge of downstream tasks—an assumption that rarely holds in open-ended, real-world scenarios.
To address this issue, another line of research explores locally deployable prompt optimizers (Lu et al., 2025; Cheng et al., 2024). These models are trained independently of any specific task, though their training data still originates from prompts generated by advanced online LLMs, implicitly treating these outputs as optimal.
Despite their differences in deployment strategies, both online and local optimization approaches rely on prompts produced by powerful online LLMs as the gold-standard optimal prompts. This design implicitly assumes downward compatibility: the notion that prompts generated by advanced LLMs can be effectively interpreted by lightweight inference models. However, our empirical analysis challenges this assumption. As shown in Fig.1(a), prompts crafted by advanced, large-scale LLMs often contain verbose, instruction-heavy content, which may overwhelm lightweight models and degrade response quality.
This finding on the limited downward compatibility of advanced LLMs raises a crucial question: Could lightweight LLMs themselves serve as effective optimization models?
Lightweight models are typically not used for prompt optimization due to their perceived limitations in handling complex optimization processes Wang et al. ; Lu et al. (2025), as illustrated in Fig.1(b). However, our research reveals a promising alternative. We find that these models can in fact serve as effective prompt optimizers when guided by explicit and interpretable prompt design merits (Fig. 1(c)). This suggests that, when equipped with clear, learnable structures, even lightweight LLMs can function as capable optimizers—offering a cost-efficient and locally deployable alternative to online methods, particularly in low-resource or privacy-sensitive settings.
To this end, we propose MePO, a Merit-Guided Prompt Optimization model. We begin by conducting a systematic empirical analysis to identify four sets of interpretable prompt merits that consistently characterize high-quality prompts, offering actionable insights into effective prompt construction. Guided by these merits, we build a prompt optimization preference dataset using optimized prompts generated by a lightweight LLM—eschewing reliance on advanced online LLMs as in prior work—and use it to train an end-to-end prompt optimization model. At the prompt level, our approach produces precise and clear merit guidance that can be generated by lightweight prompt optimization models. At the response level, our optimized prompts not only enhance the performance of similarly scaled inference models, but also demonstrate strong downward and upward compatibility, as the learned prompt merits generalize effectively across both large-scale and lightweight inference models. Besides, constructing the dataset using lightweight models greatly reduces API cost and external dependencies, enabling scalable, privacy-preserving, and locally deployable prompt optimization [^3] While we do not explicitly explore model-adaptive PO in this work, we posit that using the same model architecture for both the optimization and inference models may offer advantages due to shared internal representations..
Our main contributions are as follows:
1. We identify and formalize a set of interpretable prompt merits that contribute to both high-quality prompts and responses.
2. We construct a 40k-scale, merit-guided, prompt optimization preference dataset using lightweight LLMs while maintaining high optimization quality.
3. We propose MePO, a lightweight, locally deployable prompt optimization model trained on the merit-aligned dataset.
4. Evaluations of MePO at both the prompt and response levels show it matches or exceeds existing discrete APO methods and remains effective across inference models of varying capacities.
2 Empirical Analysis: What Merits a Good Prompt?
Prior APO frameworks commonly rely on advanced online LLMs (Guo et al., 2024; Yuksekgonul et al., 2024; Cheng et al., 2024) to generate optimized prompts. This approach is rooted in the assumption that advanced, large-scale LLMs are inherently better at crafting high-quality prompts (Pan et al., 2023). However, some studies raise concerns about the transferability of such prompts to smaller models. Wang et al. cautions that PromptAgent is designed to optimize prompts for state-of-the-art LLMs, yet these expert-level prompts often fail to transfer to smaller models such as GPT-2 or LLaMA, resulting in substantial performance degradation. Similarly, Lu et al. (2025) observes that small optimizers (13b) fail in difficult prompt optimization tasks.
We argue that these limitations are not solely due to the generation capacity of lightweight LLMs, but rather to the overly vague or underspecified optimization instructions used in prior work. For example, prompts such as “Given error feedback, generate a better prompt” (Wang et al., ), or “You are an expert in prompt optimization” (Lu et al., 2025), often lack concrete guidance. We find that when lightweight LLMs are explicitly instructed with detailed prompt patterns, they can generate effective optimized prompts despite their limited capacity (Fig. 1(c)). This observation motivates a key question: What merits a good prompt?
To enable lightweight LLMs as effective prompt optimizers, in this section, we analyze response-level and prompt-level characteristics to identify core merits that contribute to high-quality prompts.
2.1 Merit Discovery
Prior studies explore various merits that contribute to prompt effectiveness. Wei et al. (2022) proposes that including Chain-of-Thought (CoT) reasoning in prompts enhances performance, while Lampinen et al. (2022) finds that Detailed Explanations improve prompt quality. Bsharat et al. (2023) introduces 26 prompting principles, categorized into dimensions such as Prompt Structure and Clarity, and Specificity and Information. Ye et al. (2024) proposes three key components for complex reasoning prompts: Detailed Descriptions, Context Specification, and Step-by-Step Reasoning.
Although each of these works contributes valuable insights, they often emphasize different dimensions, resulting in a lack of consensus on what constitutes a high-quality prompt. To further explore this space, we take an empirical approach to uncover the merits of effective prompts through comparative evaluation.
Empirical Analysis: We randomly select 1,400 raw questions from the Alpaca dataset Taori et al. (2023) and ask a lightweight LLM [^4] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct to rewrite each question five times [^5] Implementation details are provided in Appx.B.. We then use the same LLM as the inference model to generate responses for both the raw and rewritten questions. The quality of each response is assessed using DeepSeek-R1 (Guo et al., 2025)5. Interestingly, we observe that even small changes in a prompt—sometimes just a few words—can lead to significant differences in response scores. This suggests that certain rewrites yield better prompts, which in turn lead to more effective answers.
Based on this observation, we identify raw-rewritten prompt pairs and categorize them as: (1) prompts that led to higher-scoring responses, and (2) those that resulted in lower-scoring responses, with a score difference greater than 4 points. We then ask DeepSeek-R1 to identify the merits that make the higher-scoring prompts more effective compared to the lower-scoring ones5.
Figure 2: Key merits of high-performing prompts extracted from DeepSeek-R1 evaluations.
Fig. 2 shows the top 10 merits most frequently associated with high-scoring prompts. Among them, several merits exhibit notably higher frequencies, revealing that high-quality prompts consistently reflect a set of core merits that can be effectively generated by lightweight LLMs. Motivated by this finding, we propose the following design merits to guide the construction of optimal prompts:
$\bullet$ Clarity: The optimal prompt should set clear, unambiguous expectations for the responder to enable a thorough and accurate reply.
$\bullet$ Precision: Use more precise and purposeful language, especially when referring to selecting words or concepts without a fixed pattern.
$\bullet$ Concise Chain-of-Thought: Include brief yet contextually rich reasoning or structural cues to guide the responder’s thought process, while remaining focused and concise.
$\bullet$ Preserve Original Information: Ensure no information or intent from the original prompt is lost or omitted in the transformation.
Response-Level Evaluation: To assess whether merit-aligned prompts improve response4 quality, we compare responses generated from merit-guided optimized prompts against those from raw prompts across two datasets (2,000 samples each), with scores assigned by DeepSeek-R15.
$\bullet$ Alpaca Dataset [^6] https://huggingface.co/datasets/tatsu-lab/alpaca: Raw responses are generated by text-davinci-003.
$\bullet$ BPO Dataset [^7] https://huggingface.co/datasets/THUDM/BPO: Raw responses are sourced from human-preferred answers.
Figure 3: Response-level win rate comparison. ‘Win’ indicates that the optimized prompt’s response received a higher score than that from the raw prompt.
Figure 4: Alignment Stages of Merit-Guided Prompt Optimization (MePO)
As shown in Fig. 3, prompts optimized using our designed merits—despite being applied to a lightweight model—can match or even outperform GPT-3.5-generated responses (Alpaca) and significantly beat human-preferred responses (BPO).
Prompt-Level Evaluation: To move beyond response quality and directly assess prompt quality—a dimension largely overlooked in prior work—we randomly interleaved optimized and raw prompts and asked DeepSeek-R1 to evaluate them in isolation (i.e., without seeing the generated responses)5. Across 2000 prompt comparisons per dataset, optimized prompts were preferred in 95.75% of cases on BPO and 97.9% on Alpaca. These results reinforce the effectiveness of our proposed prompt merits and demonstrate that prompt quality can be reliably evaluated independently of generated outputs.
Conclusion: Our evaluations demonstrate that optimized prompts grounded in our proposed interpretable merits not only yield stronger responses but also exhibit higher prompt quality. Moreover, the results show that effective prompt optimization is not exclusive to advanced, large-scale LLMs such as GPT-4. With merit-based guidance, lightweight LLMs can also serve as capable prompt optimizers, offering scalable and privacy-friendly alternatives for real-world deployment.
3 Method: From Merits to Optimizer
To incorporate the proposed merits into the APO pipeline, we apply them at the optimization stage as supervision signals for preference learning [^8] Alternatively, the merits could be used in conjunction with response feedback to iteratively guide the optimization process through the inference model. In this work, we focus on the optimization stage and leave integration with inference feedback for future exploration..
Our objective is to end-to-end optimize arbitrary input prompts $P_{\text{silver}}$ into refined prompts $P_{\text{golden}}$, enabling inference models to generate higher-quality responses. To achieve this, we introduce MePO, a lightweight LLM-based prompt optimization model designed for local deployment. MePO is trained under a preference learning framework, where supervision is provided by prompt constructed according to the merits identified in Sec. 2. An overview of MePO is shown in Fig. 4.
3.1 Constructing the Merit-Aligned Prompt Preference Dataset
To train MePO, we construct a Prompt Optimization Preference (POP) dataset grounded in interpretable prompt merits. Following prior work (Cheng et al., 2024; Lu et al., 2025), we adopt Alpaca and BPO as our base datasets. The instruction field in Alpaca and the prompt field in BPO are treated as raw prompts $P_{\text{silver}}$, which are refined using our defined merits to generate $P_{\text{golden}}$.
We use a base model $M_{\text{ref}}$ to generate responses $\hat{R}_{\text{golden}}$ from each optimized prompt $P_{\text{golden}}$, and retrieve the corresponding raw responses $\hat{R}_{\text{silver}}$ from Alpaca’s output and BPO’s bad_res. Each response is then evaluated by DeepSeek-R1, and we retain only those pairs that satisfy two criteria: (1) $\hat{R}_{\text{golden}}$ receives a higher evaluation score than $\hat{R}_{\text{silver}}$, and (2) the score of $\hat{R}_{\text{golden}}$ exceeds 8999According to our scoring rubric in Fig. 13, responses with scores above 8 are considered accurate and well-aligned with the prompt.. The retained samples are relabeled as $R_{\text{golden}}$ and $R_{\text{silver}}$, forming a structured 4-tuple used in training:
$(P_{\text{silver}},P_{\text{golden}},R_{\text{silver}},R_{\text{golden}}).$
Since Alpaca and BPO contain relatively well-formed prompts—either produced by text-davinci-003 or curated by humans—we further simulate real-world usage by randomly selecting 10% of $P_{\text{silver}}$ and intentionally degrading them using a base model. These lower-quality prompts better reflect common user inputs and are paired with their corresponding $P_{\text{golden}}$ to enhance data diversity.
Prompt Source Statistics Alpaca BPO
Category Naive Degraded Naive Degraded
# Prompts 25,526 3,000 10,225 1,400
Total samples: 40,151
Validation (Win Rate) GPT-4o (5k) DeepSeek-R1 (40k)
Response Win Rate 94.82% 100.00%
Prompt Win Rate 97.68% 97.62%
Table 1: Overview of our POP dataset. The top section reports prompt statistics paired from Alpaca and BPO, including 10% degraded examples. The bottom section shows validation results from DeepSeek-R1 and GPT-4o, with win rate denoting the proportion where $R_{\text{golden}}$ or $P_{\text{golden}}$ outperforms its silver counterpart.
Inference Model Optimizer ARC-Easy ARC-Challenge GSM8K BBH PiQA Avg.
Qwen2-7b - 80.68 65.19 76.88 45.42 80.20 69.67
Inference Model 81.22 66.21 79.01 48.66 81.34 71.29
BPO 80.89 66.38 77.38 50.07 82.64 71.47
FIPO 82.37 67.49 82.71 49.38 82.48 72.89
MePO 83.33 68.52 83.12 53.42 83.46 74.37
Tulu2-7b-dpo - 45.37 26.54 30.63 32.68 70.02 41.05
Inference Model 46.40 29.35 32.93 35.13 70.78 42.92
BPO 47.98 31.57 32.08 35.96 71.60 43.84
FIPO 50.55 36.95 33.66 36.06 72.63 45.97
MePO 55.05 38.14 35.18 39.89 73.60 48.37
LLaMA2-7b-chat-hf - 35.40 29.27 15.06 31.08 49.95 32.15
Inference Model 36.74 29.52 16.27 33.49 51.41 33.47
BPO 38.30 30.89 19.85 36.34 52.56 35.59
FIPO 36.24 29.44 17.03 35.81 51.58 34.02
MePO 39.86 31.74 25.25 38.68 55.33 38.17
LLaMA3-8b-instruct - 46.72 41.31 55.04 35.21 65.56 48.77
Inference Model 52.02 42.32 57.87 37.83 67.19 51.45
BPO 50.88 44.37 58.39 39.33 71.81 52.96
FIPO 49.12 44.97 56.81 39.29 75.56 53.15
MePO 55.01 51.88 59.59 42.19 76.66 57.07
Table 2: Performance comparison of prompt optimization methods across multiple datasets and inference models. ‘-’ denotes the original unoptimized prompt. ‘Inference Model’ refers to using the inference model itself as the optimizer (based on the instruction format in Fig. 11). Detailed BBH task-level results are provided in Appx.D.3.
After filtering and aggregation, the final POP dataset comprises 40k samples. Table 1 summarizes the dataset composition and presents quality validation scores from both DeepSeek-R1 and GPT-4o, showing consistently strong preference for the optimized prompts across both response- and prompt-level evaluations.
3.2 Training the Merit-Guided Prompt Optimizer
To guide the model toward generating more preferred prompts, we train the optimizer using the Direct Preference Optimization (DPO) framework (Rafailov et al., 2023). The objective is to learn a prompt optimizer that prefers merit-aligned prompts yielding higher-quality responses.
In our setup, each training instance includes a chosen prompt $P_{\text{golden}}$ that yields a higher-quality response and a rejected prompt $P_{\text{silver}}$ with a lower-quality response.
The DPO objective is formulated as:
$\mathcal{L}_{\text{DPO}}(M_{o};M_{\text{ref}})=-\mathbb{E}_{(x,\text{c},\text{r})\sim\mathcal{D}}\left[\log\sigma\left(\beta\cdot\Delta\right)\right]$
$\Delta=\log\frac{M_{o}(\text{c}\mid x)}{M_{\text{ref}}(\text{c}\mid x)}-\log\frac{M_{o}(\text{r}\mid x)}{M_{\text{ref}}(\text{r}\mid x)}$
where $M_{o}$ denotes the optimization model, and $M_{\text{ref}}$ is a reference model with adapters disabled.
Further implementation details on dataset construction and optimizer training are in Appx.C.
4 Experiments
We evaluate the effectiveness of MePO from the following perspectives: (1) Comparison with SOTA local prompt optimizers (Sec. 4.2); (2) Ablation study with training-free prompting (Sec. 4.3); (3) Upward compatibility in larger inference models (Sec. 4.4); (4) Comparison with online-based meta-prompts (Sec. 4.5) and (5) Effectiveness analysis of optimized prompts (Sec. 4.6).
4.1 Setups
MePO is trained using Qwen2.5-7B-Instruct4, selected for its strong generation ability among similarly sized models Qwen et al. (2025). This choice ensures that the optimizer is built on a lightweight LLM that maximizes performance within its size class—allowing efficient deployment without sacrificing generation quality.
Datasets We evaluate our method on five benchmark datasets: ARC-Easy, ARC-Challenge (Clark et al., 2018), GSM8K (Cobbe et al., 2021), BigBench-Hard (BBH) (Suzgun et al., 2023), and PiQA (Bisk et al., 2020). Following Lu et al. (2025), we adopt 3-shot for GSM8K, BBH, and PiQA, and zero-shot for the ARC. Prompt optimization is applied only to the test query, while the in-context examples remain unchanged.
Baselines We compare MePO with two state-of-the-art local prompt optimization models: BPO (Cheng et al., 2024) and FIPO (Lu et al., 2025). We evaluate optimizers’ performance across three inference model families: Qwen (Yang et al., 2024), Tulu (Ivison et al., 2023) and LLaMA (Touvron et al., 2023). We also use the inference model itself as an optimization baseline, motivated by the inherent alignment between a model and its own preferences, as leveraging an LLM’s natural capability to refine prompts for itself better reflects its aligned outputs (Xiang et al., 2025).
Additional implementation details and dataset descriptions are provided in Appx.D.
4.2 Main Results
We present the performance of each optimizer, including our proposed MePO, on different inference models across five datasets. The results, summarized in Table 2, lead to the following conclusions:
Efficacy MePO consistently outperforms all baselines across datasets and task models, with average accuracy gains of 1.48% on Qwen2-7B, 2.40% on Tulu2-7B, 2.58% on LLaMA2-7B, and 3.92% on LLaMA3-8B. Notable improvements are observed on ARC-Challenge with LLaMA3-8B (+6.91%), GSM8K with LLaMA2-7B (+5.4%), and ARC-Easy and BBH with Tulu2-7B (+4.5%, +3.83%).
Robustness Inference models often perform better when the optimizer shares the same architecture, reflecting alignment with model-specific preferences. For example, in the BPO–FIPO comparison, BPO (based on LLaMA2) performs better when responding with LLaMA2-7B, while FIPO (based on Tulu2) excels with Tulu2-7B. However, both optimizers show misalignment on models like LLaMA3 and Qwen2. On ARC-Easy, inference model optimization outperforms BPO (+1.14% with LLaMA3, +0.33% with Qwen2) and FIPO (+2.90% with LLaMA3). Similarly, on GSM8K, the inference model surpasses BPO (+1.63% with Qwen2) and FIPO (+1.06% with LLaMA3). In contrast, MePO consistently outperforms inference model optimization across all datasets, indicating that its learned prompt merits are broadly interpretable across architectures, demonstrating both robustness and strong generalization capability.
Inference Model Optimizer ARC-Easy ARC-Challenge GSM8K BBH PiQA Avg.
Qwen2-7b Training-Free 82.79 68.17 81.52 49.65 82.32 72.89
MePO 83.33 68.52 83.12 53.42 83.46 74.37
Tulu2-7b-dpo Training-Free 49.58 34.30 32.93 37.66 72.60 45.41
MePO 55.05 38.14 35.18 39.89 73.61 48.37
LLaMA2-7b-chat-hf Training-Free 36.91 30.55 20.91 36.60 53.10 35.61
MePO 39.86 31.74 25.25 38.68 55.33 38.17
LLaMA3-8b-instruct Training-Free 52.36 50.77 58.25 39.58 71.69 54.53
MePO 55.01 51.88 59.59 42.19 76.66 57.07
Table 3: Comparison between training-free merit-guided prompts and MePO. Detailed BBH results are in Appx.D.3.
Downward Compatibility FIPO shows performance degradation compared to the inference model optimizer on LLaMA2-7b (-0.5% on ARC-Easy, -0.08% on ARC-Challenge) and LLaMA3-8b (-1.06% on GSM8K, -2.9% on ARC-Easy). As FIPO is a 70b optimizer trained on prompts optimized by GPT-4, its outputs may be too complex for smaller inference models to interpret—indicating a lack of downward compatibility. In contrast, MePO, while built upon a relatively strong base model compared to LLaMA2-7b and LLaMA3-8b Qwen et al. (2025), maintains strong downward compatibility, consistently achieves better performance. More analyses on smaller inference models are in Appx.D.1.
4.3 Ablation Study: Training-Free Comparison
As summarized in Table 1, our filtering process retains 25.5k samples from the original 30k Alpaca entries and 10k from 14k BPO entries, discarding cases where merit-optimized prompts failed to outperform their raw counterparts. This ensures that MePO is trained only on successful optimization examples, encouraging the model to learn effective and reliable merit application.
To determine whether MePO’s performance gains result purely from the use of high-quality prompt merits or from learning how to apply them effectively, we compare it against a training-free baseline. This baseline uses our identified merits as static templates, without any learned optimization. As shown in Table 3, MePO consistently outperforms the training-free variant across all datasets and model backbones. Notably, average accuracy improves by 2.96% on Tulu2 and over 2.5% on LLaMA2 and 3, demonstrating the value of end-to-end optimization over training-free prompting.
4.4 Case Study: Upward Compatibility
To evaluate MePO’s upward compatibility, we examine whether its optimized prompts improve response quality in large-scale inference models.
LLaMA2-13b-chat-hf Tulu2-70b-dpo
- MePO - MePO
ARC-Easy 49.83 55.93 80.47 85.23
ARC-Challenge 45.05 52.73 58.87 60.49
GSM8K 28.73 35.97 61.41 64.67
BBH 36.48 43.72 48.05 49.74
PiQA 56.13 60.99 77.75 80.14
Table 4: Response evaluation of MePO prompts in larger LLMs. Detailed BBH results are in Table 9 and 14.
As shown in Table 4, applying MePO prompts to larger inference LLMs consistently improves performance across all datasets. These results indicate that MePO—despite being a lightweight prompt optimizer—generates clear and concise reasoning guidance that are well interpreted by large-scale LLMs, demonstrating strong upward compatibility.
4.5 Case Study: Meta-Prompt Comparison
To further validate MePO’s effectiveness, we compare it on GSM8K with three online-based methods: APO Pryzant et al. (2023), APE Zhou et al. (2022), and PE2 Ye et al. (2024). These methods enhance reasoning by appending an optimized meta-prompt after the query, keeping the original query unchanged while MePO directly optimizes each query. As shown in Table 5, MePO outperforms all baselines, indicating that per-query merit-driven optimization is more effective than applying a fixed meta-prompt across diverse inputs.
Optimizer Tulu2-7b-dpo LLaMA2-7b-chat-hf
APO 29.19 22.97
Iterative APE 29.72 20.92
PE2 32.15 22.37
MePO 35.18 25.25
Table 5: Comparison between meta-prompts and MePO. Details of the applied prompts are in Appx.D.2.
4.6 Case Study: Interpretability of MePO
To understand the effectiveness of MePO’s optimized prompts, we conduct a qualitative analysis on samples from the GSM8K dataset.
Example 1
Raw: Carla is downloading a 200 GB file. Normally she can download 2 GB/minute, but 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes. Then Carla has to restart the download from the beginning. How load does it take to download the file?
MePO: Carla is attempting to download a 200 GB file at a rate of 2 GB per minute. At 40% completion, her download is interrupted for 20 minutes to allow for a Windows update installation. After this interruption, she must restart the entire download from the beginning. Calculate the total time required to complete the file download, including the initial download time, the update interruption, and the subsequent restart.
Observation: The optimized prompt offers clearer structure and explicitly outlines the three components required for correct reasoning. This clarity facilitates more accurate downstream responses.
Example 2
Raw: Shiela bought five cell phones for \$150 each for a 3-month installment. A 2% interest will be charged for each unit. How much will Shiela pay each month for 3 months?
MePO: Shiela purchased five cell phones at \$150 each, with a 3-month installment plan that includes a 2% total interest rate on each phone. Calculate the total monthly payment Shiela will make over the 3-month period, considering the interest applied to each phone separately.
Observation: The raw prompt is ambiguous—it’s unclear whether the 2% interest is monthly or total. The MePO version resolves this ambiguity by clearly specifying a 2% total interest rate, enabling the model to reason correctly.
These examples illustrate how MePO’s optimized prompts reflect our defined merits—clarity, precision, and concise CoT guidance—to reduce ambiguity and improve task performance.
5 Related Work
Automatic Prompt Optimization (APO) offers a practical alternative to fine-tuning by focusing on prompt engineering instead of parameter updates. Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
Online-Based Discrete APO Several works perform discrete prompt optimization using high-capacity, API-based LLMs as optimization models. For example, Zhou et al. (2022) use InstructGPT; Wang et al. ; Guo et al. (2024); Pryzant et al. (2023) rely on GPT-3.5; Xiang et al. (2025) adopt Claude 3.5 Sonnet; and Yuksekgonul et al. (2024) utilize GPT-4o. While effective, these approaches depend on proprietary APIs, introducing practical concerns around cost, inference latency, and privacy risks.
Local Discrete APO Locally trainable prompt optimizers have been explored for general use. FIPO (Lu et al., 2025) trains a Tulu2-based model using a GPT-3.5/4-generated preference dataset. BPO (Cheng et al., 2024) uses a LLaMA-based optimizer trained on ChatGPT-optimal prompts. MAPO (Chen et al., 2023) adopts a model-adaptive strategy with GPT-3.5-generated prompts but requires a separate warm-up set per task, limiting generalization. In contrast, our model uses merit-guided prompts from a lightweight LLM, achieving strong performance without API reliance.
Prompt Merits Exploration To improve prompt quality, Arora et al. shows that prompts in open-ended QA formats outperform restrictive ones. Wei et al. (2022) introduces intermediate reasoning steps in prompts. Zhou et al. (2023) propose decomposing prompts into simpler subcomponents. Besides, several studies provide design principles for prompt optimization (Bsharat et al., 2023; Ye et al., 2024). However, these merits are largely proposed heuristically and validated through downstream experiments, leading to inconsistencies in perspective. In contrast, we derive prompt merits by prompt–response empirical analysis, grounding them in measurable improvements.
6 Conclusion
In this work, we introduce MePO, a lightweight, locally deployable prompt optimization model trained under a merit-guided preference framework. Empirical analyses are conducted to discover prompt merits—clarity, precision, and concise chain-of-thought—that contribute to high-quality prompts, resulting in a prompt preference dataset proposed using the lightweight LLMs. Experimental results show that MePO not only generates structurally clear and precise prompts, but also exhibits strong downward and upward compatibility, maintaining robust performance across variously scaled inference models. Our findings demonstrate that, with well-defined optimization merits, even lightweight LLMs can serve as effective prompt optimizers—enabling scalable, cost-efficient, and privacy-friendly deployment in real-world settings.
7 Limitations
While MePO demonstrates strong performance across diverse tasks and model scales, several limitations remain: (1) Lack of Interactive Feedback: MePO currently operates as a one-shot prompt optimizer, without incorporating iterative signals from users or inference model feedback. In real-world deployments, both user interactions and model responses can provide valuable guidance for continual prompt refinement. Integrating MePO into an interactive, feedback-driven optimization loop remains a promising direction for future work. (2) Limited Model Adaptation: Although MePO exhibits robustness across different inference models, further gains may be possible by aligning the base architectures of the optimization and inference models. While our results show that MePO performs well even under architectural mismatch, explicitly training model-adaptive optimizers could improve performance by leveraging shared internal representations.
Acknowledgements
The research is supported by the Home Team Science and Technology Agency.
Appendix A Empirical Study: Prompt Optimization Across Model Scales
Evaluating GPT-4 Optimized Prompts on Smaller Inference Models.
We adopt the optimized prompts from Lu et al. (2025), which are derived from a GPT-4 optimization dataset and used to train a Tulu2-70B-DPO model. We then evaluate the effectiveness of these prompts on smaller inference models (LLaMA2-7B and Tulu2-7B-DPO) by generating responses and scoring them with DeepSeek-R1 (see Fig. 13 for scoring setup).
LLaMA2-7B-chat-hf Response:
"Sure, here’s a traditional haiku about the ocean:
Waves roll in, peaceful
Soothing sounds, endless blue
Lullaby of sea
In this haiku, I wanted to capture the peaceful and soothing quality of ocean waves, much like a lullaby. The image of waves rolling in and out of the shore is a common one in haiku poetry, and I wanted to emphasize the sense of endlessness and tranquility that it evokes. The use of the word ‘peaceful’ in the first line sets the tone for the rest of the poem, and the image of ‘endless blue’ in the second line conveys the vastness and depth of the ocean. The final line, ‘Lullaby of sea,’ adds a sense of melody and rhythm to the poem, which is fitting given the soothing nature of ocean waves."
Tulu2-7B-DPO Response:
"Haiku is a traditional form of Japanese poetry that consists of three lines with a syllable pattern of 5-7-5. It is a concise and evocative form that aims to capture a moment or emotion in a few words. A lullaby, on the other hand, is a song that is typically sung to put children to sleep. It often has a soothing and calming melody, and the lyrics are usually simple and repetitive.Here’s my traditional haiku about the ocean:
Gentle ocean waves
Crash upon the shore in peace
Lullaby of the sea
In this haiku, I’ve captured the peaceful and soothing quality of ocean waves, as they gently crash upon the shore. The repetition of the "ocean waves" phrase in the first and third lines creates a sense of continuity, emphasizing the endless cycle of the ocean’s"
These examples show that prompts crafted by powerful optimization models may not be directly interpretable by smaller models due to verbosity or overly rich CoT structure, leading to inconsistent outputs.
Evaluating Lightweight LLMs with Complex Optimization Algorithms.
We assess whether lightweight LLMs can follow optimization prompts designed for advanced models. Specifically, we use the Differential Evolution (DE) algorithm from Guo et al. (2024), which guides prompt generation using EvoPrompt instructions. We supply the EvoPrompt Prompt and DE-generated Optimal Prompt as input to the model and use DeepSeek-R1 to compare their quality, as shown in Fig. 5.
Evaluating Merits-Guided Prompt Optimization with Lightweight LLMs.
We further test whether lightweight LLMs can act as effective optimization models when guided by clear, interpretable merits. Using the Basic Prompt from Guo et al. (2024), we apply our proposed merit-based instruction (see Fig. 14) to generate optimized prompts. We then evaluate the difference between Basic and Optimized Prompts using DeepSeek-R1 with the pairwise setup shown in Fig. 9.
These studies collectively show that lightweight LLMs struggle with vague or complex optimization instructions but can perform competitively when provided with structured, merit-aligned guidance.
Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
To explore the characteristics of effective prompts, we conduct a comparative evaluation using a series of specially designed prompts.
Prompt Rewriting.
We begin by using a rewriting prompt (shown in Fig. 6) to generate multiple reformulations of each raw question. Specifically, five rewritten versions are produced for each input prompt to preserve diversity. Below is an example from the Alpaca dataset, where the raw prompt results in an incorrect response, while Rewrite 1 and Rewrite 2 yield correct answers under the same inference model4:
Raw: What does DNA stand for?
Rewrite 1: Can you tell me what the acronym DNA represents?
Rewrite 2: What is the full form of DNA?
Prompt Effectiveness Evaluation.
To evaluate prompt quality at the prompt level (i.e., independent of response content), we design a comparison prompt that asks DeepSeek-R1 to assess which of the two prompt variants is better and explain why, as illustrated in Fig. 8. The summarized features provided by DeepSeek-R1 are then collected for merit discovery. Below presents two examples from DeepSeek’s evaluations, highlighting which prompt was preferred and why, along with the associated merit types.
Example 1
DeepSeek Evaluation:
###Explicit Guidance: The Golden Prompt provides a clearer direction by specifying the focus on "the importance of customer service," which helps the responder structure their thoughts around a central theme rather than a vague topic.
###Contextual Richness: By emphasizing "the importance," the Golden Prompt encourages the responder to include reasons, examples, or impacts, leading to a more substantive and engaging tweet.
###Engagement Potential: The Golden Prompt is more likely to generate a tweet that resonates with readers, as discussing "importance" naturally invites the responder to highlight value or benefits, making the content more compelling and shareable.
In contrast, the Sliver Prompt ("Craft a tweet regarding customer service experiences") is too broad and could lead to generic or unfocused responses, such as simple complaints or praises without deeper insight.
Example 2
DeepSeek Evaluation:
###Precision in Terminology: The Golden Prompt uses the more accurate linguistic term "form" instead of "conjugation," which is more appropriate for discussing verb structures in English grammar.
###Clarity of Focus: By placing the verb phrase in quotation marks ("will have helped"), it clearly identifies the exact linguistic element to be analyzed, reducing ambiguity.
###Technical Accuracy: The phrasing aligns better with grammatical analysis, as "form" encompasses tense, aspect, and mood, which are all relevant to "will have helped" (future perfect tense).
###Conciseness: It removes the unnecessary framing of "Identify the conjugation of the verb:" and gets straight to the point, making it more efficient for analysis.
Prompt Preference Judgement.
We use a prompt template (see Fig. 9) to conduct direct pairwise comparisons between two prompts, with judgments provided by DeepSeek-R1. The positions of the prompts are randomly assigned to mitigate positional bias.
Response Scoring.
The quality of the model’s response to each prompt is evaluated using DeepSeek-R1. The scoring template is displayed in Fig. 13, inspired by Yang et al. (2025).
Prompt Optimization.
The prompt template used to refine raw prompts into merit-guided optimized versions is shown in Fig. 14.
Appendix C POP Dataset Construction and Training Details
Prompt Degradation.
To simulate noisy user inputs—such as “how can i go fr sigapor?”—we degrade 10% of the raw prompts from the Alpaca and BPO datasets using a base model. The prompt used for degradation is shown in Fig. 7. Below is an example from the Alpaca dataset, generated by the model referenced in 4:
Raw: Describe the atmosphere at the beach.
Degraded: describ the atmossphre at the bheach.
DPO Input Construction.
To train MePO under the DPO objective, we formulate each training instance using the format shown in Fig. 10, where S_P is the silver (unoptimized) prompt $P_{\text{silver}}$, S_R is its corresponding response $R_{\text{silver}}$, and G_R is the preferred response $R_{\text{golden}}$.
Training Configuration.
We fine-tune MePO based on Qwen2.5-7B-Instruct4 using 4 NVIDIA RTX A5000 GPUs. Each input instruction $x$ is truncated or padded to 2000 tokens. Training hyperparameters and cost are summarized in Table 6.
Qwen2.5-7b-instruct
Node1 1
Per_GPU_batch 1
Accumulations 4
HyperParams
Epoch=2, Seq Len=2048,
lr=1e-6, beta=0.01,
Top P=0.95,
Temperature=0.8,
loss type=’sigmoid’,
Train (2 epoch) 1.5 day
Table 6: Training setup and hyperparameters used for fine-tuning MePO.
Appendix D Dataset Information and Experimental Implementation
We evaluate MePO on four multiple-choice and one generative benchmark dataset:
Evaluation Protocol.
GSM8K, PiQA, and BBH are evaluated in a 3-shot setting to facilitate answer extraction, while all other datasets are evaluated in a zero-shot setting. For consistency, all BBH tasks are reformatted into multiple-choice format, except for multistep_arithmetic_two, object_counting, and word_sorting, which retain their original formats. PiQA, which includes two-solution questions, is also converted to multiple-choice format. For all multiple-choice datasets, we determine the answer by selecting the option (e.g., A, B or 1, 2) with the highest next-token logit probability following the ##Answer: token.
Prompt Templates.
Fig. 12 shows the prompt formats used for GSM8K, PiQA and ARC. For BBH, the multiple-choice tasks use the PiQA prompt while the math tasks use the GSM8K prompt.
Baseline Implementations.
We compare MePO with two state-of-the-art prompt optimizers. The prompt instructions used for inference models serving as prompt optimizers are shown in Fig. 11. Details of the baselines are as follows:
All optimization models, including MePO, are loaded in 8-bit precision for prompt generation. Inference models are run in 4-bit mode with a generation length below 512 to reduce memory overhead.
Inference Model Optimizer ARC-Easy ARC-Challenge GSM8K BBH PiQA Avg.
LLaMA3.2-1b-instruct - 27.02 23.13 16.38 31.05 49.56 29.43
Inference Model 28.79 24.23 20.24 32.29 49.73 31.06
BPO 27.99 24.04 15.85 31.77 49.89 29.91
FIPO 25.19 23.55 21.30 31.72 50.27 30.41
MePO 29.42 24.40 21.46 33.40 50.44 31.82
LLaMA3.2-3b-instruct - 55.46 73.34 63.23 35.6 66.32 58.79
Inference Model 60.75 78.16 63.00 37.91 69.37 61.84
BPO 55.74 74.75 56.71 37.34 67.14 58.34
FIPO 67.66 78.66 65.13 37.65 69.53 63.73
MePO 61.43 79.58 65.66 40.49 70.29 63.49
Table 7: Downward compatibility: Response evaluation of optimizer in lightweight inference models.
D.1 Case Study: Downward Compatibility
To further analyze the downward compatibility of optimization models, we evaluate performance using two smaller inference models: LLaMA-3.2-3B-Instruct [^16] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and LLaMA-3.2-1B-Instruct [^17] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct.
As shown in Table 7, both BPO and FIPO exhibit performance degradation when paired with lightweight inference models. For example, under LLaMA-3.2-3B-Instruct, BPO drops by 5.01% on ARC-Easy, 3.41% on ARC-Challenge, and 2.23% on PiQA. Under LLaMA-3.2-1B-Instruct, FIPO drops by 3.60% on ARC-Easy, and BPO drops by 4.39% on GSM8K. Both also show performance decreases on BBH under both lightweight inference models. In contrast, MePO consistently improves performance across all datasets, demonstrating strong downward compatibility and robustness in low-resource settings.
D.2 Case Study: Online-Based Discrete APO Comparison
To further validate MePO’s effectiveness in optimizing prompts or instructions with refined meta-prompts, we compare MePO with three discrete APO methods:
  • APO, a prompt optimization method that performs optimization via textual gradients using GPT-3.5-Turbo-Instruct in a text-based dialogue setting;
  • Iterative APE, a template-based strategy in which GPT-3.5 generates a pool of candidate prompts, followed by selection based on development set performance;
  • PE2, a prompt engineering method that optimizes prompts through three key components using GPT-3.5-Turbo-Instruct.
The meta-prompts for each method are listed in Table 8, adapted from Ye et al. (2024), and are prepended before the ##Answer segment shown in Fig. 12(a).
Optimizer Meta Prompt
APO
Given the scenario, perform necessary
calculations and provide a step-by-step
explanation to arrive at the correct
numerical answer. Consider all
information provided.
Iterative APE
Let’s dissect this and tackle it gradually,
one phase at a time.
PE2
Let’s solve the problem step-by-step and
calculate the required total value correctly.
Table 8: Meta-prompts used for each baseline discrete APO method.
D.3 BBH Experimental Results
Detailed results for the 23 BBH tasks are presented in the following tables: Table 9 (raw prompts), Table 10 (optimized by the inference model), Table 11 (training-free merit-guided optimization), Table 12 (BPO), Table 13 (FIPO), and Table 14 (MePO).
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA2-13b
-chat-hf
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
Tulu2-70b
-dpo
date_understanding 44.8 30.8 28.8 31.6 42 25.6 27.2 62
disambiguation_qa 63.2 32.8 31.2 32.4 30 36.4 33.6 68.8
hyperbaton 75.6 48.4 48.4 48.4 48.4 51.6 48.4 78
logical_deduction_five_objects 49.2 22.4 18 28 33.2 24.4 19.6 50
logical_deduction_seven_objects 42.4 17.6 14.8 30 27.2 30.4 15.2 42.8
logical_deduction_three_objects 66.4 32.8 31.6 33.6 36.4 35.2 31.6 71.2
movie_recommendation 48 26.8 22.4 33.2 23.6 23.2 22.4 52.8
penguins_in_a_table 54.11 21.23 22.6 31.51 46.58 26.71 23.97 50.68
reasoning_about_colored_objects 47.6 32.4 32 16.4 29.2 30 16.4 46
ruin_names 37.2 24 30.8 35.2 27.2 41.6 28.4 66.8
salient_translation_error_detection 32 26.8 23.2 21.6 28 34 16.8 52
snarks 60.11 44.94 46.07 46.07 50.56 57.3 46.07 81.46
word_sorting 6 3.2 2.8 3.2 4.8 3.2 1.2 10.4
temporal_sequences 61.2 19.2 28.4 28.4 33.2 28 28.4 62
tracking_shuffled_objects_five_objects 28.4 19.2 20 20.8 30.8 20 20 17.2
tracking_shuffled_objects_seven_objects 13.2 24 14.4 13.6 22 11.2 12.8 13.6
tracking_shuffled_objects_three_objects 27.2 29.2 24.4 26.8 49.2 30 30.8 28
causal_judgement 62.57 49.2 51.87 51.34 49.2 47.59 51.87 53.48
formal_fallacies 54 52.8 53.2 53.6 52.4 51.6 53.2 52
navigate 57.6 53.6 42 41.6 41.6 41.2 42 52.8
web_of_lies 46.4 48.8 48.8 49.2 54 44.4 46.4 46.8
sports_understanding 70 50 46.4 70.4 54 53.6 43.2 45.6
boolean_expressions 34.8 56.8 52.4 54.8 51.2 48 54 30.4
multistep_arithmetic_two 33.6 12.4 9.6 36 14 46 16 16.8
object_counting 36 33.2 32.8 50.4 33.2 48.8 46.8 49.6
Table 9: Results of raw BBH dataset across the evaluated inference models.
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
date_understanding 54.4 34 34 32 27.6 31.2
disambiguation_qa 62.8 34 32.4 34.4 40.4 36.8
hyperbaton 77.6 49.2 49.2 49.2 51.6 48.4
logical_deduction_five_objects 50.4 22.8 19.2 36.4 29.2 20.8
logical_deduction_seven_objects 44 17.2 15.6 31.2 29.6 16.4
logical_deduction_three_objects 74.8 34.8 32 38.8 40 32
movie_recommendation 51.6 30.4 23.6 35.2 24.4 26.8
penguins_in_a_table 56.16 26.03 23.29 32.19 28.77 23.29
reasoning_about_colored_objects 48.4 35.2 34 18.8 28.4 18
ruin_names 38.4 26.4 32.4 38 43.6 28.4
salient_translation_error_detection 38 21.2 24.8 23.6 35.2 20.4
snarks 65.73 48.88 49.44 47.19 57.3 46.07
word_sorting 10.4 5.6 4.8 5.2 4.8 1.2
temporal_sequences 62.8 31.2 33.2 30.4 28.4 28.4
tracking_shuffled_objects_five_objects 29.6 29.6 30.8 32.4 20 22.4
tracking_shuffled_objects_seven_objects 15.2 25.6 23.2 14.4 14 13.6
tracking_shuffled_objects_three_objects 28 30 26 28.4 31.2 32
causal_judgement 72.73 52.41 53.2 50.8 55.61 52.41
formal_fallacies 54.8 54.4 53.2 51.6 52 53.2
navigate 60.8 56 43.2 42.8 41.6 42
web_of_lies 47.2 48.8 48.8 50 48.8 47.6
sports_understanding 75.6 63.45 48 72 57.6 45.6
boolean_expressions 41.6 57.2 53.6 52.4 47.6 54
multistep_arithmetic_two 36.4 10.8 12 37.2 54 18
object_counting 39.2 33.2 37.2 58.4 56 48.4
Table 10: Results of the BBH dataset optimized by the inference model across the evaluated inference models.
- Qwen2-7b Tulu2-7b-dpo LLaMA2-7b-chat-hf LLaMA3-8b-instruct
date_understanding 55.6 36 34 48
disambiguation_qa 66 37.2 35.2 35.2
hyperbaton 78.4 50 48.4 49.6
logical_deduction_five_objects 55.6 25.2 19.2 34.8
logical_deduction_seven_objects 44 28.4 21.6 38.8
logical_deduction_three_objects 75.2 34.4 42.4 40.4
movie_recommendation 53.2 31.2 24.4 37.6
penguins_in_a_table 56.16 27.4 23.29 36.99
reasoning_about_colored_objects 50.8 34 35.2 20.8
ruin_names 38 26.4 40 39.6
salient_translation_error_detection 38.4 34 37.2 18.8
snarks 62.36 50 50.56 48.88
word_sorting 13.6 6.8 5.6 5.2
temporal_sequences 64.4 36.8 36.8 31.2
tracking_shuffled_objects_five_objects 29.6 38.4 34.8 33.2
tracking_shuffled_objects_seven_objects 14.4 24.8 23.2 14.4
tracking_shuffled_objects_three_objects 30 31.6 31.6 31.2
causal_judgement 55.08 52.41 51.87 52.41
formal_fallacies 54 54 54 52
navigate 57.2 58.8 43.2 42.4
web_of_lies 51.2 49.6 49.2 51.6
sports_understanding 76 64.8 49.2 79.6
boolean_expressions 41.6 60.4 53.6 54.4
multistep_arithmetic_two 41.2 12.6 12 41.6
object_counting 39.2 36.4 58.4 50.8
Table 11: Results of the BBH dataset optimized by training-free merit-guidance across the evaluated inference models.
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
date_understanding 56.8 33.2 35.2 48.8 29.6 29.2
disambiguation_qa 63.2 34 34.4 40.4 44.4 34.4
hyperbaton 80 53.2 50 49.6 51.6 48.8
logical_deduction_five_objects 52.8 25.2 24 31.2 29.2 20.4
logical_deduction_seven_objects 46.4 28.8 19.2 35.6 24.8 15.2
logical_deduction_three_objects 78.4 32.8 45.6 39.2 34.8 32
movie_recommendation 61.2 22.8 24.4 37.6 23.2 26.8
penguins_in_a_table 54.11 23.97 23.29 36.3 29.45 23.97
reasoning_about_colored_objects 52.4 34 35.6 20.8 28.4 18
ruin_names 38.4 28 36.8 35.6 38 28.4
salient_translation_error_detection 38.4 27.6 35.2 21.2 33.6 18.8
snarks 66.85 53.37 53.37 48.88 53.93 46.07
word_sorting 12.4 5.6 6 6 5.2 3.2
temporal_sequences 66.4 26.4 38.4 30.8 35.6 28.4
tracking_shuffled_objects_five_objects 29.6 29.2 31.2 34 20 19.6
tracking_shuffled_objects_seven_objects 14.4 28 23.2 15.2 14 13.2
tracking_shuffled_objects_three_objects 27.6 31.6 31.6 31.6 31.6 32
causal_judgement 53.48 50.8 52.04 52.94 52.94 51.87
formal_fallacies 52 53.2 54.14 54 52.4 53.2
navigate 59.6 58.4 44.8 48.4 44.8 42
web_of_lies 49.2 51.6 48.8 49.6 49.2 48.8
sports_understanding 76.4 61.2 53.2 74.8 55.2 45.6
boolean_expressions 45.6 63.6 55.6 53.2 50.4 54
multistep_arithmetic_two 36.8 10 17.6 37.2 46.8 16.8
object_counting 34.4 32.4 34.8 50 54.4 43.6
Table 12: Results of the BBH dataset optimized by BPO across the evaluated inference models.
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
date_understanding 55.2 38.4 29.2 36 30 27.2
disambiguation_qa 61.6 34.8 31.2 40.4 32 33.2
hyperbaton 78.8 33.2 48.4 60 52 48.4
logical_deduction_five_objects 50 21.6 22.8 34.4 23.6 20.4
logical_deduction_seven_objects 46.4 25.2 20.4 39.2 23.2 15.6
logical_deduction_three_objects 70.8 30.4 42 39.2 38.8 32
movie_recommendation 62 36.8 23.97 31.6 34.4 24.4
penguins_in_a_table 54.11 36.3 23.29 30.82 52.05 23.29
reasoning_about_colored_objects 50.8 31.6 33.2 24.8 25.6 16.4
ruin_names 46 34.4 36 38.8 40 29.2
salient_translation_error_detection 37.6 32.4 33.6 24.4 37.2 18.8
snarks 62.92 46.07 50.56 46.07 55.62 46.07
word_sorting 8.4 4.8 3.2 7.6 4.4 4.8
temporal_sequences 63.2 27.6 28.4 28.8 40 28.8
tracking_shuffled_objects_five_objects 31.2 24.8 32.4 36.8 20.8 20.8
tracking_shuffled_objects_seven_objects 16.8 25.6 24.8 14.4 14.4 16.4
tracking_shuffled_objects_three_objects 31.2 30.4 31.6 30.4 31.6 30.8
causal_judgement 62.57 53.48 51.87 51.34 55.08 52.41
formal_fallacies 53.2 54.4 54 52.4 53.2 53.2
navigate 59.2 60.8 42 46.4 53.2 42
web_of_lies 58.8 48 53.2 45.2 49.2 48.8
sports_understanding 81.2 66.4 47.26 71.6 55.6 46
boolean_expressions 39.6 58.8 53.6 61.2 39.6 54.4
multistep_arithmetic_two 16.4 11.6 17.2 39.6 24.4 16.4
object_counting 36.4 33.6 61.2 50.8 55.2 43.2
Table 13: Results of the BBH dataset optimized by FIPO across the evaluated inference models.
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA2-13b
-chat-hf
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
Tulu2-70b
-dpo
date_understanding 57.77 36.4 35.6 49.2 52.4 34 31.2 50
disambiguation_qa 66.4 38.4 37.2 48.8 54.8 44.8 40.8 68.8
hyperbaton 81.6 54.4 50.4 50.4 54.4 56.8 48.4 62
logical_deduction_five_objects 58.8 27.2 24.4 47.2 42.4 30 23.6 52
logical_deduction_seven_objects 49.2 29.6 20.4 39.6 36.8 21.2 18.4 51.6
logical_deduction_three_objects 82 36.8 51.2 42.8 39.2 40.4 34.8 65.6
movie_recommendation 63.2 38.4 26.8 38.4 29.2 32 27.2 34
penguins_in_a_table 56.85 28.77 25.34 38.36 54.79 31.51 25.34 51.37
reasoning_about_colored_objects 56.8 35.6 38 23.6 41.6 26 18.4 58.8
ruin_names 38 27.6 41.6 42 28.8 45.6 28.4 73.6
salient_translation_error_detection 43.2 36.8 38.8 28.8 31.2 39.2 23.6 54.8
snarks 67.42 53.37 51.12 49.44 66.85 60.11 46.07 83.15
word_sorting 16.8 8.4 8 8 10.7 7.3 3.2 16.8
temporal_sequences 38 37.6 39.2 33.2 42 29.2 28.8 66.8
tracking_shuffled_objects_five_objects 31.2 39.6 33.2 35.2 32.8 20.4 23.2 21.2
tracking_shuffled_objects_seven_objects 15.2 30 26.4 16.4 28.4 18.8 15.2 18
tracking_shuffled_objects_three_objects 33.6 32.4 33.2 33.2 64.8 31.6 31.6 29.6
causal_judgement 49.2 56.68 51.87 51.79 56.15 62.57 52.41 59.36
formal_fallacies 51.2 57.2 54.8 53.2 52.8 53.2 53.2 59.2
navigate 61.6 58.8 44.4 44.8 48.4 47.2 42 60
web_of_lies 51.2 55.2 51.6 51.6 59.2 51.2 49.6 49.6
sports_understanding 77.6 67.2 54.8 77.2 61.2 58 46 49.2
boolean_expressions 48.4 61.2 54 54.4 54 52.8 55.6 36.4
multistep_arithmetic_two 45.6 14.4 14.4 41.2 16 57.6 18.4 18
object_counting 39.6 35.2 60.4 55.2 34 60.8 49.6 53.6
Table 14: Results of the BBH dataset optimized by MePO across the evaluated inference models.
Appendix E Prompt Templates
The prompts used in this work are listed below:
Figure 5: Prompt used to evaluate EvoPrompt Prompt and Optimal Prompts generated by the EvoPrompt algorithm under lightweight LLMs.
Figure 6: Prompt used to rewrite raw questions.
Figure 7: Prompt used to generate degraded prompts. S_P denotes the original raw prompt.
Figure 8: Prompt used to evaluate the effectiveness of two prompts. S_P denotes the prompt yielding a lower-scoring response; G_P yields a higher-scoring response.
Figure 9: Prompt used to compare raw and optimized prompts. The two prompts are randomly placed in S_P and G_P to mitigate position bias.
Figure 10: Prompt used to construct DPO training inputs. S_P, S_R, and G_R denote $P_{\text{silver}}$, $R_{\text{silver}}$, and $R_{\text{golden}}$, respectively.
Figure 11: Prompt used to optimize prompts for inference model. S_P denotes the original raw prompt.
(a) GSM8K
(b) PiQA
(c) ARC
Figure 12: Prompt formats used for downstream task evaluation.
Figure 13: Prompt used by DeepSeek-R1 to score responses. {question} is replaced by the prompt; {modelresponse} is replaced by the model’s response.
Figure 14: Prompt for Prompt Optimization. S_P denotes the prompt that needs to be optimized.